# From Freshness to Effectiveness: Goal-Oriented Sampling for Remote Decision Making

### v2

@@ VIII-C Discussions and Insights
Moreover, Fig. 7 demonstrates that the performance gap between the proposed goal-oriented sampling policy and the myopic decision-making policy gradually narrows as the average delay $\mathbb{E}[Y_{i}]$ increases. This convergence reflects a fundamental limitation in decision-making under communication constraints: when the delay becomes excessively large, the received state information is so outdated that it loses its relevance to the current system condition. As a result, the optimal goal-oriented decision making degenerates into state-independent behavior when the average delay is large. In such regimes, the value of state information diminishes, and the benefits of optimized information transmission vanish.
Furthermore, Fig. 7 demonstrates that when the average delay is relatively small, AoI-optimal, zero-wait, and constant-wait sampling strategies also perform competitively, approaching the performance of goal-oriented sampling. However, their effectiveness deteriorates rapidly under large delays, sometimes falling below that of the constant-action myopic policy. This occurs because when delays are minimal, the received state information is still timely enough to inform near-optimal decisions. Conversely, under large average delay, the received data becomes too stale to offer meaningful advantage for decision making.
@@
In contrast, the uniform sampling policy in Fig. 8 exhibits a U-shaped cost curve. As $f_{\max}$ decreases, its performance converges to that of AoI-optimal sampling policy. This convergence highlights a key insight: under sparse sampling constraints, the transmitted information becomes too outdated/uninformative to support effective state-dependent decision-making, resulting in uniformly poor performance. Conversely, when the sampling interval $d=1/f_{\max}$, the system begins to accumulate a queuing backlog, resulting in increased delivery delays. This queuing-induced staleness reduces the timeliness of information at the decision-maker, thereby degrading goal-oriented decision-making performance.
IX Conclusion

### v3

@@ VIII-C Discussions
A key observation from Fig. 7 is that the myopic decision-making policy, which selects a fixed action regardless of state information, yields a constant cost independent of delay. This myopic policy serves as a delay-agnostic baseline in the sense that it does not explicitly condition its decision rule on AoI beyond the last received state; the gap to the proposed goal-oriented sampling quantifies the benefit of explicitly leveraging AoI in decision-making. This highlights an insight: in scenarios where the control policy is insensitive to informed state information, enhancing the communication channel (e.g., by reducing delay) may not lead to any improved goal-oriented decision-making performance. Moreover, Fig. 7 demonstrates that the performance gap between the proposed goal-oriented sampling policy and the myopic decision-making policy gradually narrows as the average delay $\mathbb{E}[Y_{i}]$ increases. This convergence reflects a fundamental limitation in decision-making under communication constraints: when the delay becomes excessively large, the received state information becomes so outdated that it no longer provides reliable guidance for current decisions. As a result, the optimal goal-oriented policy degenerates into state-independent behavior in high-delay regimes, where the influence of timely and fresh information on decision performance becomes negligible.
Fig. 9 illustrates the average cost achieved by various sampling strategies under a truncated geometric delay model with parameter $q=0.3$. The $x$-axis represents the cutoff parameter $Y_{\max}$, which limits the maximum possible delivery delay. As $Y_{\max}$ increases, the truncated delay distribution gradually approaches the standard (non-truncated) geometric distribution, and the resulting average cost curves converge accordingly. Across the entire range of $Y_{\max}$, the proposed goal-oriented sampling strategy consistently outperforms all baselines, achieving significantly lower average cost—especially under large delay cutoffs.
@@
In contrast, the uniform sampling policy in Fig. 10 exhibits a U-shaped cost curve. As $f_{\max}$ decreases, its performance converges to that of AoI-optimal sampling policy. This convergence highlights a key insight: under sparse sampling constraints, the transmitted information becomes too outdated to support effective state-dependent decision-making, resulting in uniformly poor performance. Conversely, when the sampling interval $d=1/f_{\max}$ becomes small, the system begins to accumulate a queuing backlog, resulting in increased delivery delays. This queuing-induced staleness reduces the timeliness of information at the decision-maker, thereby degrading goal-oriented decision-making performance.
IX Conclusion and Future Directions